OpenStudy (anonymous):

Nathan had an infection, and his doctor wanted him to take penicillin. Because Nathan’s father and paternal grandfather were allergic to penicillin, Nathan has a 75% chance of having the same allergy. The doctor performed a skin test to see whether Nathan would react to it. The test is 98% accurate. If the test predicts that Nathan is not allergic to penicillin, what is the probability after the test that Nathan is not allergic to the drug? A. 0.1135 B. 0.2450 C. 0.9423 D. 0.9932

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

you know how to do this? @Koikkara

OpenStudy (koikkara):

@sugahead12 Nathan has 0.75 chance of allergy 0.25 chance of no allergy test is 0.98 correct 0.02 incorrect Nathan is not allergic to penicillin and the test predicts it or in other words Nathan no allergy and test incorrect multiply the probabilities: Pr(no allergy) * Pr(incorrect) = 0.25*0.02 = 0.005

OpenStudy (anonymous):

Yeah i get that but how do i get the real answer after multiplying that

OpenStudy (koikkara):

@sugahead12 Seems like it's been answered daily, have a look at closed questions too ! Nice to meet You !! Have A Good Day !!

OpenStudy (anonymous):

Hey if you still want the answer it's C; 0.9423